Hello there,
I'm having issues with word problems.
The sum of two numbers is 17 and their product is 42. Find the two numbers.

Answer by scott8148(6628)

"The sum of two numbers is 17" ___ x + y = 17 ___ x = 17 - y

"their product is 42" ___ x*y = 42

substituting ___ x(17-x) = 42 ___ 17x - x^2 = 42

0 = x^2 - 17x + 42

factoring ___ 0 = (x - 14)(x - 3)

x is 3 or 14 and y is the other one